A 143-g sample of mercury is at an initial temperature of 25 °C. If 1067 joules of heat are applied to the sample, what is the final temperature of the mercury? The specific heat capacity of mercury is 0.14 J/(g∙ °C).
Answer #1

Following is the - complete Answer -&- Explanation: for the given Question, in.....typed format....

Answer:

Final temperature of the ( sample ) mercury : Tf = 78.30 oC   

Explanation:

Following is the complete Explanation, for the above Answer, in.....typed format...

  • Given:
  1. ​​​​​​​ Mass of the sample, of mercury: m = 143.0 g ( grams )
  2. Initial Temperature of the sample; mercury : Ti = 25.0 oC
  3. Amount of Heat applied to the sample: Q = 1067 J ( Joules )
  4. Specific Heat Capacity of mercury : C = 0.14 J / (g . oC )
  • ​​​​​​​Step - 1:

​​​​​​​ Let the final temperature , of the sample , -- be : TfoC   

We know the following: Equation of Heat Transfer :

Q = m x C x T   ------------------------------------------Equation - 1

Where:

  1. m = mass of the sample ( i.e. mercury )
  2. C = specific heat capacity of the sample ; [=] J / ( g. oC )
  3. T = Change in temperature , of the sample, = ( Tf - Ti ) = ( Tf - 25.0 oC ) , i.e. in the given case.
  4. Q =  Amount of Heat transferred, in the process of heat transfer....
  • ​​​​​​​Step - 2:

​​​​​​​Plugging in the values in Equation - 1, as above, we would get the followiong:

Q = m x C x T   ------------------------------------------Equation - 1

  ( 1067 J ) = ( 143.0 g ) x ( 0.14 J / ( g. oC ) ) x (  Tf- 25.0 oC )

  (  Tf- 25.0 oC ) =   ( 1067 J ) / [  ( 143.0 g ) x ( 0.14 J / ( g. oC ) ) ] =  53.296 53.30  oC

Tf = ( 53.30 + 25.0 ) =  78.30 oC   .....

Final Temperature : Tf = 78.30 oC   

​​​​​​​